The present age of suraj and deeraj is in the ratio 3:4. Sic years hence,the ratio of 1/3th of the age of suraj to 1/5th of the age of deeraj will be 4:3. What is the present age of suraj?

Grade:12th pass

1 Answers

Shubham Pandey
6 years ago
Let present age of Suraj be 3x and that deeraj be 4x, so that ratio of their ages is 3:4After six yrs., suraj`s age will be 3x+6 and deeraj`s age will be 4x+6According to question:1/3(3x+6) / 1/5(4x+6) = 4/3On simplifying, x=6So present age of Suraj= 3x = 18 yrs